# Algebraic Topology Principles behind Topological Quantum Error Correction

### v1

@@ III.2 Intersections on 2-Manifolds and TQEC Ability
The original toric code [21, 23] can be used as quantum memory because the ground state of the Hamiltonian in equation 8 has a four-fold degeneracy. Therefore, no local stabilizer checks can tell apart the degenerate ground states. The four ground states can be transformed into one another by applying logical $\sigma_{X}$ and $\sigma_{Z}$ operations, which anti-commute because the chain of logical $\sigma_{X}$ and $\sigma_{Z}$ operators intersect at an odd number of sites.
Thus, we propose that the TQEC ability of any manifold $M$ is related to the intersection properties of sub-manifolds of $M$. The definition of homotopy of curves and transverse intersection is presented in Appendix A.1, and we shall only consider transverse intersection points. Here, the formal definition of the modulo-2 intersection is presented, which is useful for proving the theorems in later sections [34, 26, 27].
@@
[…]
Where ##\## counts the number of intersections. If $f$ is the inclusion of $A$ as a sub-manifold, then we may use the notation $I_{2}(A,B):=I_{2}(f,B)$.
@@
The proof can be found in any differential topology textbook, for instance [34], and the mathematical definition of homotopy is in Appendix A.3. Intuitively, a homotopy of a curve means a smooth and continuous deformation of that curve. The theorem III.1 is crucial when we investigate the necessary conditions for a topological manifold to be used for TQEC purposes. Restricting to 2-manifolds, the following corollary is obtained.
@@
Given a closed and compact 2-manifold $M$. For any closed loops $A,B\subset M$, and $\partial A=\partial B=0$. Then, the parity of the number of intersection points between loop $A$ and $B$ is invariant under homotopies of $A$ and $B$.
@@ III.3 TQEC on 2-Dimensional Manifolds
In this paper, we considered the “loop-based TQEC”, where logical $\sigma_{X}$ is defined by operating $\sigma_{X}$ on any cycles $\gamma\in C_{1}(M;\mathbb{Z}_{2})$, such that $\gamma\subset E$, where $\partial[\gamma]=0$ and $[\gamma]\neq 0$; a logical $\sigma_{Z}$ is defined to operate on any co-cycles, $\beta\in C^{1}(M;\mathbb{Z}_{2})$, such that $\beta\subset E$, where $\delta[\beta]=0$ and $[\beta]\neq 0$, just as we defined in section III.1. Then, we have the following theorem.
@@
A 2-dimensional closed and compact manifold can be used as TQEC codes for qubits if and only if there exist at least two closed loops with an odd number of intersections.
Proof.
Firstly, using the basic principle of QEC, by proper stabilizer measurements, the set of all possible quantum errors will collapse into a set of discrete quantum errors. Thus, for any possible qubit error $Err$, it can be written as $Err=e_{0}I+e_{1}\sigma_{X}+e_{2}\sigma_{Z}+e_{3}\sigma_{X}\sigma_{Z}$ for some constant $e_{0},e_{1},e_{2},e_{3}$. This means, for QEC purposes, a method to implement logical $\sigma_{X}$ and $\sigma_{Z}$ on the code is sufficient.
With the theorem III.2, and applying basic algebraic topology properties of surfaces, the first trivial observation is that all simply connected surfaces cannot be used for TQEC purposes. We formalize it using the following definition and a corollary [26].
@@
Proof.
Theorem III.2 states that, for any closed surfaces to be used as TQEC code, we require at least 2 loops that interact in an odd number of points. And theorem III.1 states that the modulo-2 intersection number is homotopic invariant. However, for any 2 non-overlapping loops on simply connected surfaces, each can be shrunk down to a point continuously, where they do not intersect. This means that two non-overlapping loops can only transversely intersect at an even number of points. Thus, they cannot be used as TQEC codes. ∎
As a result, simply connected surfaces, such as spheres ($S^{2}$), cannot be used for TQEC purposes. This may seem rather trivial, but more theorems are building up from it.
@@
Proof.
Notation-wise, since $\mathbb{Z}_{2}$ is communtative, we will represent it as the modulo-2 addition group that the elements of $\mathbb{Z}_{2}$ are $\{0,1\}$, and for simplicity, for $a\in\{0,1\}$ we represent $\sum_{i=1}^{n}a:=a^{n}$.
Then, to prove the “if” part, considering elements belonging to the $H_{1}(M;\mathbb{Z}_{2})$, $\gamma\in C_{1}(M;\mathbb{Z}_{2}),[\gamma]\neq 0\text{ and }\partial\gamma=0$. And elements belongs to the $H^{1}(M;\mathbb{Z}_{2})$, $\beta\in C^{1}(M;\mathbb{Z}_{2}),[\beta]\neq 0\text{ and }\delta\beta=0$. The Poincaré duality as in the theorem II.1 states that $H_{1}(M;\mathbb{Z}_{2})\cong H^{1}(M;\mathbb{Z}_{2})$. Thus, the pairing given by a function $p:H_{1}(M;\mathbb{Z}_{2})\times H^{1}(M;\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}$ is well-defined, and the $\mathbb{Z}_{2}$ coefficients ensure that $[\gamma]^{2}=0$ and $[\beta]^{2}=0$. By the properties of elements belonging to the first cohomology group, for arbitrary integers $r,s$, we have $[\beta]^{2r+1}([\gamma]^{2s+1})=[\beta]([\gamma])=1$, while $[\beta]^{2r}([\gamma]^{s})=[\beta]^{r}([\gamma]^{2s})=0$. The Poincaré duality states that since $H_{1}(M;\mathbb{Z}_{2})$ is isomorphic to $H^{1}(M;\mathbb{Z}_{2})$, thus, there is a bijective mapping between elements in $C^{1}$ to elements in $C_{1}$, so since $[\gamma]$ is a class of closed loops, $[\beta]$ are also closed loops on the surface $M$. Therefore, logical $\sigma_{X}$ can be defined on $[\gamma]$ and logical $\sigma_{Z}$ can be defined on $[\beta]$, where they anti-commute, which defines a TQEC code.
We can visualize the effect of the co-homology cycle on the homology cycle using the Toric code as an example, for a torus $T^{2}$. In Figure 1, where the blue cycle corresponds to $\gamma\in C_{1}(T^{2};\mathbb{Z}_{2}),\partial\gamma=0,[\gamma]\neq 0$, and the red cycle corresponds to $\beta\in C^{1}(T^{2};\mathbb{Z}_{2}),\delta\beta=0,[\beta]\neq 0$. Since both $\gamma$ and $\beta$ are non-trivial, they intersect at an odd number of sites, and the parity of the intersection site is invariant under curve homotopy. Therefore, the torus can be used as TQEC codes for qubits.
Theorem III.3 characterizes the logical qubit encoding capabilities of different 2-manifolds. The 2-manifolds can be classified into two primary categories. The first consists of orientable surfaces $M_{g}$ of genus $g$, constructed iteratively via the connected sum operation with a torus, i.e., $M_{g}=M_{g-1}\#T^{2}$, see Appendix A.5 for a formal definition of connected sum. By convention, $M_{g}$ is equivalently denoted as $gT^{2}:=(g-1)T^{2}\#T^{2}$. The second category comprises non-orientable surfaces, denoted as $gP^{2}$, defined recursively through connected sums with the real projective plane: $gP:=(g-1)P\#\mathbb{R}P^{2}$ for $g\geq 1$ [24]. Another famous theorem in Algebraic Topology, that has been discussed widely [35, 24], is quoted below.
@@
    The surface $gP^{2}$: The first homology group of the surface $H_{1}(gP;\mathbb{Z}_{2})=\mathbb{Z}_{2}^{g}$. Thus, by theorem III.3, it can encode $g$ qubits. Also, when $g=2$, the surface $\mathbb{R}P^{2}\#\mathbb{R}P^{2}\cong K$
Remark that the 2-dimensional real projective plane $\mathbb{R}\text{P}^{2}$, the Klein bottle $K$, and their generalization $gP^{2}$ surface are all non-orientable, and they cannot be used for TQEC codes for qudits with dimension $D>2$ as proven in [24]. If we recall that definition of TQEC codes as in section III.1, it is defined on 2-dimensional cell complexes, which can be represented as a graph, $G=(V,E,F)$. In the discussion above, we only considered the cell complexes that are embedded on 2-manifolds, because of the properties of manifolds. However, in theory, TQEC codes can be defined on cell complexes not embedded in manifolds. The problem is, if a cell complex $X$ is not embeddable on any manifold, then a lot of nice properties of manifolds, like the Poincaré duality as in theorem II.1, do not hold anymore. We will briefly discuss the formation of TQEC on general 2-dimensional cell complexes in Appendix B.
@@ IV TQEC Codes on Higher Dimensional Manifolds
We start the section by presenting the construction of TQEC codes on higher-dimensional manifolds. The idea is that when generalizing the TQEC from 2-manifolds to an $n$-manifold $M$, the code is defined upon a cell complex $X_{M}$ embedded on the manifold $M$, where a physical qubit is placed on each $i$-cell $D^{i}_{\alpha}$ for $11\leq i\leq n-$ 1, where $\alpha$ denotes the indices, and logical qubits are encoded separately on each dimension.
Logical $\sigma_{X}$ is operated on homology cycles $\gamma\in C_{i}(M;\mathbb{Z}_{2})$, that $\partial\gamma=0,[\gamma]\neq 0$. And logical $\sigma_{Z}$ is operated on cohomology cycles, $\beta\in C^{i}(M;\mathbb{Z}_{2})$, that $\delta\gamma=0,[\gamma]\neq 0$. As shown in equation 7, if $i+j=n$, then the pairing between $H_{i}(M,\mathbb{Z}_{2})$ and $H_{j}(M,\mathbb{Z}_{2})$ over $\mathbb{Z}_{2}$ is well-defined, meaning that it is possible to find $n$-manifold $M$ with $n>2$, and for some $11\leq i\leq n-$ 1 we can define logical $\sigma_{X}$ and $\sigma_{Z}$ operations that their modulo-2 intersections is well-defined, and as from the Poincaré duality II.1, $H^{i}(M,\mathbb{Z}_{2})\cong H_{n-i}(M,\mathbb{Z}_{2})$. Here, we will develop a general theory of TQEC using an $n$-dimensional cell complex embedded on an $n$-manifold.
@@
A cell complex $X_{M}$ embedded on a $n$-dimensional closed and compact manifold $M$ can encode qubits on the $i$-th dimensional skeleton $X^{i}$, where $11\leq i\leq n-$ 1, if and only if the homology group $H_{i}(M;\mathbb{Z}_{2})$ is non-trivial. The number of qubits that can be encoded on the $i$-cycles equals the rank $H_{i}(M;\mathbb{Z}_{2})$, or equivalently, the $i$-th Betti number $b_{i}(M;\mathbb{Z}_{2})$.
The definition of the $n$-th skeleton of the cell complex $X_{M}$ is formally defined in Appendix A.2. Remark: The reason why we only consider the $H_{i}(M)$ for $11\leq i\leq n-$ 1 is that all closed and path-connected manifold having $H_{0}(M;\mathbb{Z}_{2})\approx\mathbb{Z}_{2}$, and the 0-simplex is essentially a collection of points. By definition, points have no boundary [26, 27], but boundary operations are considered when defining the stabilizer checks in TQEC. Since we cannot formulate boundary maps and stabilizer checks on points, considering $H_{0}(M)$ and $H_{n}(M)$ is not useful for TQEC.
Proof.
Firstly, we prove that the closed and compact $n$-manifold $M$ can be used for TQEC code for qubits on the $i$-cycle and $i$-cocycle only if its $i$-th homology group is nontrivial. The reason being that if $H_{i}(M;\mathbb{Z}_{2})=0$, then $b_{i}(M;\mathbb{Z}_{2})=0$. There is only 1 element in the group $H_{i}(M;\mathbb{Z}_{2})$, which means all closed $i$-cycles are equivalently null-homologous. This implies all $\gamma\in C_{i}(M;\mathbb{Z}_{2}),\partial\gamma=0$ are equivalent, that $\gamma^{s}=\gamma$ for all positive integer $s$. By the definition of elements belonging to the cohomology group $H^{i}(M;\mathbb{Z}_{2})$, even if there exist $\beta\in C^{i}(M;\mathbb{Z}_{2}),[\beta]\neq 0\text{ and }\delta\beta=0$. For all non-negative integer $r$, $\beta(\gamma^{r})=\beta(\gamma)=0$. However, for $\sigma_{X}$ operator, $\sigma_{X}^{2r+1}=\sigma_{X}$ and $\sigma_{X}^{2r}=\mathbb{I}$ for an arbitrary integer $r$. This implies that $\sigma_{X}$ cannot be defined on any $i$-cycles; hence, defining a TQEC code is impossible.
Then, to prove the “if” part, considering elements belonging to the $H_{i}(M;\mathbb{Z}_{2})$, $\gamma\in C_{i}(M;\mathbb{Z}_{2}),[\gamma]\neq 0\text{ and }\partial\gamma=0$. And elements belongs to the $H^{i}(M;\mathbb{Z}_{2})$, $\beta\in C^{i}(M;\mathbb{Z}_{2}),[\beta]\neq 0\text{ and }\delta\beta=0$. The Poincaré duality as in the theorem II.1 states that $H_{i}(M;\mathbb{Z}_{2})\cong H^{n-i}(M;\mathbb{Z}_{2})$. Thus, the pairing given by a function $p:H_{i}(M;\mathbb{Z}_{2})\times H^{i}(M;\mathbb{Z}_{2})\rightarrow\mathbb{Z}_{2}$ is well-defined. The $\mathbb{Z}_{2}$ coefficients ensure that $[\gamma]^{2}=[\beta]^{2}=0$. And by the definition of elements belonging to the cohomology group, for arbitrary integers $r,s$, we have $[\beta]^{2r+1}([\gamma]^{2s+1})=[\beta]([\gamma])=1$, while $[\beta]^{2r}([\gamma]^{s})=[\beta]^{r}([\gamma]^{2s})=0$. $[\gamma]$ represent the class of non-trivial closed $i$-cycles and $[\beta]$ represent the class of non-trivial closed $i$-cocycles, which is isomorphic to the class of non-trivial closed $(n-i)$-cycles. Therefore, logical $\sigma_{X}$ can be defined on $[\gamma]$ and logical $\sigma_{Z}$ can be defined on $[\beta]$, where they anti-commute, which defines a TQEC code.
@@ IV.1 Examples of higher-dimensional TQEC: The 3-Torus Code
Figure 3: Examples of possible closed loops $\gamma^{1}$ in the 3-torus $T^{3}$, where $\gamma^{1}\in C_{1}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{1}=0$. (a) $[\gamma^{1}]=0$, which is a contractible loop. (b) $[\gamma^{1}]\neq 0$, which is a non-contractible loop.
Figure 4: Examples of possible closed 2-torus $\gamma^{2}$ in the 3-torus $T^{3}$, where $\gamma^{2}\in C_{2}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{2}=0$. (a) $[\gamma^{2}]=0$, which is a compressible 2-torus. (b) $[\gamma^{2}]\neq 0$, which is a non-compressible 2-torus.
Firstly, for $i=\{1,2\}$, we would like to provide some examples of $\gamma^{i}\in C_{i}(T^{3};\mathbb{Z}_{2})$ and $\partial\gamma^{i}=0$ for both $[\gamma^{i}]=0$ and $[\gamma^{i}]\neq 0$. Figure 3 shows examples of contractible and non-contractible loops in $T^{3}$, $\gamma^{1}\in C_{1}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{1}=0$, for both $[\gamma^{1}]=0$ and $[\gamma^{1}]\neq 0$. And Figure 4 shows examples of compressible and non-compressible 2-torus $T^{2}$ in $T^{3}$, $\gamma^{2}\in C_{2}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{2}=0$, for both $[\gamma^{2}]=0$ and $[\gamma^{2}]\neq 0$. For simplicity, denote $\{\gamma^{j}_{0}\}:=\{\gamma^{j}:[\gamma^{j}]=0\}$, and $\{\gamma_{1}^{j}\}:=\{\gamma^{j}\}\setminus\{\gamma^{j}_{0}\}$.
It is relatively trivial to see that $\gamma_{0}^{1}$ and $\gamma_{0}^{2}$ will always intersect at an even number of sites. Furthermore, the number of intersection sites of $\gamma_{i}^{1}$ with $\gamma_{0}^{2}$ and $\gamma_{0}^{1}$ with $\gamma_{i}^{2}$ is always even. The only exception is the number of intersections between $\gamma_{1}^{1}$ with $\gamma_{1}^{2}$, which can be odd.

### v2

@@ III.2 Intersections on 2-Manifolds and TQEC Ability
The original toric code [21, 23] can be used as quantum memory because the ground state of the Hamiltonian in equation 19 has a four-fold degeneracy. Therefore, no local stabilizer checks can tell apart the degenerate ground states. The four ground states can be transformed into one another by applying logical $\sigma_{X}$ and $\sigma_{Z}$ operations, which anti-commute because the chain of logical $\sigma_{X}$ and $\sigma_{Z}$ operators intersects at an odd number of sites.
Thus, we propose that the TQEC ability of any manifold $M$ is related to the intersection properties of sub-manifolds of $M$. The definition of homotopy of curves and transverse intersection is presented in Appendix B.1, and we shall only consider transverse intersection points. Here, the formal definition of the modulo-2 intersection is presented, which is useful for proving the theorems in later sections [34, 26, 27].
@@
Let $A$ be a compact manifold, $M$ be a manifold, and $B\subset M$ be a submanifold, all of them are without boundary and satisfying $dim(A)+dim(B)=dim(M)$. Let $f:A\rightarrow M$ be a smooth map, then the mod 2mod\ 2 intersection number $I_{2}(f,B)$ of $f$ with $B$ is defined as,
$I_{2}(f,B):=\#f^{-1}(B)\ (mod\ 2)$ (20)
Where #\# counts the number of intersections. If $f$ is the inclusion of $A$ as a sub-manifold, then we may use the notation $I_{2}(A,B):=I_{2}(f,B)$.
@@
The proof can be found in any differential topology textbook, for instance [34], and the mathematical definition of homotopy is in Appendix B.4. Intuitively, a homotopy of a curve means a smooth and continuous deformation of that curve. The theorem III.1 is crucial when we investigate the necessary conditions for a topological manifold to be used for TQEC purposes. Restricting to 2-manifolds, the following corollary is obtained.
@@
Given a closed and compact 2-manifold $M$. For any closed loops $A,B\subset M$, and $\partial A=\partial B=0$. Then, the parity of the number of intersection points between loops $A$ and $B$ is invariant under homotopies of $A$ and $B$.
@@ III.3 TQEC on 2-Dimensional Manifolds
In this paper, we considered the “loop-based TQEC”, where logical $\sigma_{X}$ is defined by operating $\sigma_{X}$ on any cycles $\gamma\in C_{1}(M;\mathbb{Z}_{2})$, such that $\gamma\subset E$, where $\partial[\gamma]=0$ and $\gamma]\neq 0$; a logical $\sigma_{Z}$ is defined to operate on any co-cycles, $\beta\in C^{1}(M;\mathbb{Z}_{2})$, such that $\beta\subset E$, where $\delta[\beta]=0$ and $\beta]\neq 0$, just as we defined in section III.1. Then, we have the following theorem.
@@
A 2-dimensional closed and compact manifold can be used as a TQEC code for qubits if and only if there exist at least two closed loops with an odd number of intersections.
The proof is in Appendix A. With the theorem III.2, and applying basic algebraic topology properties of surfaces, the first trivial observation is that all simply connected surfaces cannot be used for TQEC purposes. We formalize it using the following definition and a corollary [26].
@@
The proof is demonstrated in Appendix A. As a result, simply connected surfaces, such as spheres ($S^{2}$), cannot be used for TQEC purposes. This may seem rather trivial, but more theorems are building up from it.
@@
The proof of this theorem is in Appendix A. We can visualize the effect of the co-homology cycle on the homology cycle using the Toric code as an example, for a torus $T^{2}$. In Figure 2, where the blue cycle corresponds to $\gamma\in C_{1}(T^{2};\mathbb{Z}_{2}),\partial\gamma=0,[\gamma]\neq 0$, and the red cycle corresponds to $\beta\in C^{1}(T^{2};\mathbb{Z}_{2}),\delta\beta=0,[\beta]\neq 0$. Since both $\gamma$ and $\beta$ are non-trivial, they intersect at an odd number of sites, and the parity of the intersection site is invariant under curve homotopy. Therefore, the torus can be used as a TQEC code for qubits.
Theorem III.3 characterizes the logical qubit encoding capabilities of different 2-manifolds. The 2-manifolds can be classified into two primary categories. The first consists of orientable surfaces $M_{g}$ of genus $g$, constructed iteratively via the connected sum operation with a torus, i.e., $M_{g}=M_{g-1}\#T^{2}$, see Appendix B.6 for a formal definition of connected sum. By convention, $M_{g}$ is equivalently denoted as $gT^{2}:=(g-1)T^{2}\#T^{2}$. The second category comprises non-orientable surfaces, denoted as $gP^{2}$, defined recursively through connected sums with the real projective plane: $gP:=(g-1)P\#\mathbb{R}P^{2}$ for $g\geq 1$ [24]. Another famous theorem in Algebraic Topology, which has been discussed widely [35, 24], is quoted below.
@@
    The surface $gP^{2}$: The first homology group of the surface $H_{1}(gP;\mathbb{Z}_{2})=\mathbb{Z}_{2}^{g}$. Thus, by theorem III.3, it can encode $g$ qubits. Also, when $g=2$, the surface $\mathbb{R}P^{2}\#\mathbb{R}P^{2}\cong K$ .
Remark that the 2-dimensional real projective plane $\mathbb{R}\text{P}^{2}$, the Klein bottle $K$, and their generalization $gP^{2}$ surface are all non-orientable, and they cannot be used for TQEC codes for qudits with dimension $D>2$ as proven in [24]. If we recall the definition of TQEC codes as in section III.1, it is defined on 2-dimensional cell complexes, which can be represented as a graph, $G=(V,E,F)$. In the discussion above, we only considered the cell complexes that are embedded in 2-manifolds, because of the properties of manifolds. However, in theory, TQEC codes can be defined on cell complexes not embedded in manifolds. The problem is, if a cell complex $X$ is not embeddable on any manifold, then a lot of nice properties of manifolds, like the Poincaré duality as in theorem II.1, do not hold anymore. We will discuss the formation of TQEC on general 2-dimensional cell complexes in Section VI.
There has been previous discussion on topological codes on non-orientable manifolds, including $\mathbb{R}P^{2}$ and Klein bottles $K$. Ref. [36] studies toric, surfaces, and the honeycomb Floquet codes on non-orientable surfaces with cross‑cap defects and shows that the emergent anyon‑exchange symmetry between $e$ type and $m$ type particles, which can be implemented as a constant‑depth local logical gate. Ref. [37, 38] discussed higher-dimensional TQEC codes, for instance, the $\mathbb{Z}_{2}$ homological codes on $K\times S^{1}$. The discussion is from a more condensed matter flavor. Later, we will demonstrate the TQEC ability on higher-dimensional manifolds and the intersection properties.
@@ IV TQEC Codes on Higher Dimensional Manifolds
We start the section by presenting the construction of TQEC codes on higher-dimensional manifolds. The idea is that when generalizing the TQEC from 2-manifolds to an $n$-manifold $M$, the code is defined upon a cell complex $X_{M}$ embedded on the manifold $M$, where a physical qubit is placed on each $i$-cell $D^{i}_{\alpha}$ for $1\leq i\leq n-1$, where $\alpha$ denotes the indices, and logical qubits are encoded separately on each dimension.
Logical $\sigma_{X}$ is operated on homology cycles $\gamma\in C_{i}(M;\mathbb{Z}_{2})$, that $\partial\gamma=0,[\gamma]\neq 0$. And logical $\sigma_{Z}$ is operated on cohomology cycles, $\beta\in C^{i}(M;\mathbb{Z}_{2})$, that $\delta\gamma=0,[\gamma]\neq 0$. As shown in equation 14, if $i+j=n$, then the pairing between $H_{i}(M,\mathbb{Z}_{2})$ and $H_{j}(M,\mathbb{Z}_{2})$ over $\mathbb{Z}_{2}$ is well-defined, meaning that it is possible to find $n$-manifold $M$ with $n>2$, and for some $1\leq i\leq n-1$ we can define logical $\sigma_{X}$ and $\sigma_{Z}$ operations that their modulo-2 intersections is well-defined, and as from the Poincaré duality II.1, $H^{i}(M,\mathbb{Z}_{2})\cong H_{n-i}(M,\mathbb{Z}_{2})$. Here, we will develop a general theory of TQEC using an $n$-dimensional cell complex embedded on an $n$-manifold.
@@
A cell complex $X_{M}$ embedded on an $n$-dimensional closed and compact manifold $M$ can encode qubits on the $i$-th dimensional skeleton $X^{i}$, where $1\leq i\leq n-1$, if and only if the homology group $H_{i}(M;\mathbb{Z}_{2})$ is non-trivial. The number of qubits that can be encoded on the $i$-cycles equals the rank $H_{i}(M;\mathbb{Z}_{2})$, or equivalently, the $i$-th Betti number $b_{i}(M;\mathbb{Z}_{2})$.
The definition of the $n$-th skeleton of the cell complex $X_{M}$ is formally defined in Appendix B.2. Remark: The reason why we only consider the $H_{i}(M)$ for $1\leq i\leq n-1$ is that all closed and path-connected manifolds having $H_{0}(M;\mathbb{Z}_{2})\approx\mathbb{Z}_{2}$, and the 0-simplex is essentially a collection of points. By definition, points have no boundary [26, 27], but boundary operations are considered when defining the stabilizer checks in TQEC. Since we cannot formulate boundary maps and stabilizer checks on points, considering $H_{0}(M)$ and $H_{n}(M)$ is not useful for TQEC. This theorem is proven in Appendix A.
@@ IV.1 Examples of higher-dimensional TQEC: The 3-Torus Code
Figure 4: Examples of possible closed loops $\gamma^{1}$ in the 3-torus $T^{3}$, where $\gamma^{1}\in C_{1}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{1}=0$. (a) $\gamma^{1}]=0$, which is a contractible loop. (b) $\gamma^{1}]\neq 0$, which is a non-contractible loop.
Figure 5: Examples of possible closed 2-torus $\gamma^{2}$ in the 3-torus $T^{3}$, where $\gamma^{2}\in C_{2}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{2}=0$. (a) $\gamma^{2}]=0$, which is a compressible 2-torus. (b) $\gamma^{2}]\neq 0$, which is a non-compressible 2-torus.
Firstly, for $i=\{1,2\}$, we would like to provide some examples of $\gamma^{i}\in C_{i}(T^{3};\mathbb{Z}_{2})$ and $\partial\gamma^{i}=0$ for both $\gamma^{i}]=0$ and $\gamma^{i}]\neq 0$. Figure 4 shows examples of contractible and non-contractible loops in $T^{3}$, $\gamma^{1}\in C_{1}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{1}=0$, for both $\gamma^{1}]=0$ and $\gamma^{1}]\neq 0$. And Figure 5 shows examples of compressible and non-compressible 2-torus $T^{2}$ in $T^{3}$, $\gamma^{2}\in C_{2}(T^{3};\mathbb{Z}_{2}),\partial\gamma^{2}=0$, for both $\gamma^{2}]=0$ and $\gamma^{2}]\neq 0$. For simplicity, denote $\{\gamma^{j}_{0}\}:=\{\gamma^{j}:[\gamma^{j}]=0\}$, and $\{\gamma_{1}^{j}\}:=\{\gamma^{j}\}\setminus\{\gamma^{j}_{0}\}$.
It is relatively trivial to see that $\gamma_{0}^{1}$ and $\gamma_{0}^{2}$ will always intersect at an even number of sites. Furthermore, the number of intersection sites of $\gamma_{i}^{1}$ with $\gamma_{0}^{2}$ and $\gamma_{0}^{1}$ with $\gamma_{i}^{2}$ is always even. The only exception is the number of intersections between $\gamma_{1}^{1}$ and $\gamma_{1}^{2}$, which can be odd.